OpenStudy (anonymous):

[8.03] Factor x2 − 4x + 5 Prime (x + 5)(x − 1) (x − 5)(x − 1) (x + 5)(x + 1)

OpenStudy (anonymous):

I need one that would give me the negive 4x

OpenStudy (uri):

2 − 4x + 5 Find a number when adding or subtracting is 4 and multiplying is 5 SO, -5+1=-4 5 x 1 =5

OpenStudy (gorv):

option u gave ........in that answer is not there

OpenStudy (gorv):

x2-4x-5 this equation should be there according to ur option

OpenStudy (gorv):

@uri -5x-x=-6x

OpenStudy (uri):

-5x+1x=-4x* @gorv

OpenStudy (gorv):

@uri but u did -5x-x

OpenStudy (gorv):

we cant factor above equation ....acc to given opton

OpenStudy (gorv):

root will be imaginary

OpenStudy (gorv):

2+i and 2-i are the roots of given equation

OpenStudy (gorv):

@Sammy224

OpenStudy (anonymous):

thank you